A soccer player on a horizontal playground kicks a ball away with a velocity

v0 under an angle α with the playground. The ball can be considered to be a projectile. Air resistance can be neglected and the gravitational acceleration is g. During the trajectory, the ball reaches a maximum height H.Now he kicks a second ball under the same angle but with double velocity. What is the maximum height the second ball will reach?



1
Expert's answer
2021-11-21T17:23:59-0500

"\text{for first ball:}"

"v_y = v_0\sin\alpha"

"h_{max}= \frac{v_y^2}{2g}= \frac{(v_0\sin \alpha)^2}{2g}=H"

"\text{for second ball:}"

"v'_0= 2v_0"

"v'_y = v'_0\sin\alpha= 2v_0\sin\alpha"

"h'_{max}= \frac{(v'_0\sin \alpha)^2}{2g}= \frac{(2v_0\sin \alpha)^2}{2g}=4H"


"\text{Answer: }4H"